# Domain decomposition-based coupling of Operator Inference Reduced Order Models via the Schwarz alternating method

### v2

@@ Abstract
This paper presents and evaluates an approach for coupling together subdomain-local reduced order models (ROMs) constructed via non-intrusive operator inference (OpInf) with each other and with subdomain-local full order models (FOMs), following a domain decomposition of the spatial geometry on which a given partial differential equation (PDE) is posed. Joining subdomain-local models is accomplished using the overlapping Schwarz alternating method, a minimally-intrusive multiscale coupling technique that works by transforming a monolithic problem into a sequence of subdomain-local problems, which communicate through transmission boundary conditions imposed on the subdomain interfaces. After formulating the overlapping Schwarz alternating method for OpInf ROMs, termed OpInf-Schwarz, we evaluate the method’s accuracy and efficiency on several test cases involving the heat equation in two spatial dimensions. We demonstrate that the method is capable of coupling together arbitrary combinations of OpInf ROMs and FOMs, and that speed-ups over a monolithic FOM are possible when performing OpInf ROM coupling.
@@ 1 Introduction
Despite advancements in both computer architectures and algorithms, the modeling and simulation of complex physical systems often requires tremendous computational resources. These requirements may preclude many-query analyses such as engineering design or uncertainty quantification. While projection-based reduced order models (ROMs) have promised to mitigate this difficulty, traditional intrusive ROMs, e.g., Galerkin [13, 37] and least squares Petrov-Galerkin (LSPG) projection ROMs [4], have their own shortcomings, including a lack of systematic refinement mechanisms, a lack of robustness, stability, and accuracy in the predictive regime, and lengthy implementation time requirements.
@@
This paper presents a promising approach for mitigating the aforementioned difficulties by enabling domain decomposition- (DD-)based coupling of subdomain-local ROMs with each other and/or with subdomain-local full order models (FOMs). Our approach is based on the following ingredients: (i) a decomposition of the physical domain of interest into two or more overlapping subdomains, (ii) the construction of subdomain-local ROMs and/or FOMs in each of the subdomains, and (iii) the rigorous coupling of the subdomain-local models via the Schwarz alternating method [33]. The Schwarz alternating method is based on the simple idea that, if the solution to a partial differential equation (PDE) is known in two or more regularly shaped domains comprising a more complex domain, these local solutions can be used to iteratively build a solution for the more complex domain, with information propagating between the subdomains through carefully constructed transmission boundary conditions (BCs). We choose the Schwarz alternating method since it has a number of advantages over competing multiscale coupling methods. These advantages of the Schwarz alternating method include its concurrent nature, its ability to couple non-conformal meshes with different element topologies and different time integrators with different time steps for dynamic problems all without introducing non-physical artifacts into the solution, and its non-intrusive implementation into existing codes [23, 24].
@@
Building on our past work in developing the Schwarz alternating method as a means to couple together FOMs [23, 24], intrusive projection-based ROMs [3] and physics-informed neural networks (PINNs) [38], we focus our attention herein on advancing the method to work with a non-intrusive model order reduction (MOR) technique known as operator inference (OpInf) [26, 36]. Unlike traditional intrusive MOR, which requires access to the underlying FOM code in order to project the governing PDE(s) onto a reduced subspace, OpInf works by assuming a functional form (usually linear or quadratic [10]) for the ROM in terms of to-be-learned reduced operators, and solving an optimization problem offline for these operators. This procedure significantly reduces both the development time and the time-to-impact. While DD-based couplings between ROMs and FOMs are not new, the majority of the literature on this topic has focused on developing intrusive coupling methods (e.g., Lagrange multipliers, optimization-based coupling) used to couple intrusive ROMs; the interested reader is referred to [3, 20, 5, 14, 7, 8, 27] and the references therein for more details. Related past work on data-driven couplings using Schwarz-like methods has focused on intrusive ROMs [6, 28, 17, 16], or on utilizing the coupling to accelerate NN training [38, 18, 19]. The proposed approach is most similar to the recent work by Farcas et al. [9], which develops a DD-based coupling of subdomain-local OpInf ROMs by learning appropriate reduced operators responsible for the coupling. Since each subdomain problem is solved once in [9], rather than by performing an iteration to convergence as done within our Schwarz framework, the subdomain-local solutions must be extended to the full domain and smoothly combined to achieve a continuous solution. Due to this, the resulting extended solutions can exhibit noticeable errors near the subdomain boundaries.
@@
The remainder of this paper is organized as follows. In Section 2, we describe the overlapping version of the Schwarz alternating method applied to our model problem, the two-dimensional (2D) unsteady heat equation. In Section 2.2, we present some OpInf preliminaries. In Section 3, we describe our software implementations of the proposed Schwarz-based coupling approach applied to OpInf ROMs, hereafter called the OpInf-Schwarz method, which makes use of the open source FEniCSx [2] and OpInf [21] libraries. Results are presented in Section 4. We end with a concluding summary and a discussion of future work in Section 5.
@@ 2 The Schwarz Alternating Method Applied to Operator Inference ROMs
In the following sections, we will consider the specific model problem of the 2D heat equation, towards addressing the challenges that are encountered in this novel combination of the Schwarz method and operator inference. We stress that neither of these techniques inherently require an assumption of linearity, noting that the authors of the original operator inference paper [26] directly contrasted their method with the linearity assumption of Dynamic Mode Decomposition (DMD) [32]. While non-linearity will likely present further challenges to overcome for the OpInf-Schwarz coupling method, the heat equation provides an initial unsteady test problem that can suggest the feasibility of the method for more complicated problems.
@@ 2.1 Schwarz Alternating Method Preliminaries
$\begin{array}[]{rcrl}\dot{u}(x,t)-\Delta u(x,t)&=&f(x),&\text{in }\Omega\times[0,T],\\ u(x,t)&=&g(x),&\text{on }\partial\Omega\times[0,T]\\ u(x,0)&=&v(x),&\text{in }\Omega,\end{array}$ (1)
where $\Omega\in\mathbb{R}^{p}$ is an open bounded domain for $p=1,2,3$ with boundary $\partial\Omega$, $f(x)$ and $g(x)$ are given source and boundary condition functions, respectively, $v(x)$ defines the initial condition, and $T>0$. Suppose we decompose the domain $\Omega$ into two overlapping subdomains $\Omega_{1}$ and $\Omega_{2}$, such that $\Omega_{1}\cup\Omega_{2}=\Omega$ and $\Omega_{1}\cap\Omega_{2}\neq\emptyset$ and, for later use, define $\overline{\Omega_{1}}$ as the closure of $\Omega_{1}$ and similarly for $\Omega_{2}$. Suppose also that we decompose the time interval $[0,T]$ into a set of non-overlapping time intervals $I_{n}=[t_{n},t_{n+1}]$, where $T\geq t_{n+1}>t_{n}\geq 0$. The overlapping Schwarz algorithm solves the following sequence of subdomain-local problems:
$\left\{\begin{array}[]{rcll}\dot{u}_{1}^{(k+1)}-\Delta u_{1}^{(k+1)}&=&f^{(k+1)},&\text{in }\Omega_{1}\times[t_{n},t_{n+1}]\\ u_{1}^{(k+1)}&=&g,&\text{on }(\partial\Omega\cap\overline{\Omega_{1}})\times[t_{n},t_{n+1}]\\ u_{1}^{(k+1)}&=&u_{2}^{(k)},&\text{on }\Gamma_{1}\times[t_{n},t_{n+1}],\end{array}\right.$ (2)
@@
$\left\{\begin{array}[]{rcll}\dot{u}_{2}^{(k+1)}-\Delta u_{2}^{(k+1)}&=&f^{(k+1)},&\text{in }\Omega_{2}\times[t_{n},t_{n+1}]\\ u_{2}^{(k+1)}&=&g,&\text{on }(\partial\Omega\cap\overline{\Omega_{2}})\times[t_{n},t_{n+1}]\\ u_{2}^{(k+1)}&=&u_{1}^{(k+1)},&\text{on }\Gamma_{2}\times[t_{n},t_{n+1}],\end{array}\right.$ (3)
for Schwarz iteration $k=0,1,2,...$ subject to initial conditions $u_{1}(x,0)=v|_{\Omega_{1}}$ and $u_{2}(x,0)=v|_{\Omega_{2}}$. In (2), $u_{i}$ for $i=1,2$ denotes the solution in subdomain $\Omega_{i}$, and $\Gamma_{i}$ is the so-called Schwarz boundary (see Figure 1). It is common to set $u_{1}^{(0)}=v|_{\partial\Omega_{1}}$ and $u_{2}^{(0)}=v|_{\partial\Omega_{2}}$ when initializing the Schwarz iteration process, to ensure solution compatibility with the initial condition. A key advantage of the Schwarz alternating method is that it allows for the subdomains $\Omega_{1}$ and $\Omega_{2}$ to be discretized using different meshes and/or element types [23, 24]. In the case where $\Omega_{1}$ and $\Omega_{2}$ are discretized using different meshes and do not have a coincident interface, applying the Schwarz boundary condition on the Schwarz boundaries $\Gamma_{i}$ requires the construction of a projection operator; this can be done via a simple application of finite element interpolation functions readily available in most codes [24, 22]. The iterative process in (2) and (3) continues until a set of pre-determined criteria are met. In the present work, convergence criteria are based on the Euclidian norm of the solution differences between consecutive Schwarz iterations; that is, Schwarz is deemed converged when $\epsilon_{\text{abs}}^{(k)}<\delta_{\text{abs}}$ and $\epsilon_{\text{rel}}^{(k)}<\delta_{\text{rel}}$ for some pre-specified Schwarz tolerances $\delta_{\text{abs}},\delta_{\text{rel}}>0$, where
@@
for Schwarz iteration $k=1,2,...$. The Schwarz iteration process (2)–(3) is converged within each time interval $[t_{n},t_{n+1}]$ before moving on to the next time interval. A key advantage of this time-stepping approach is that different time integrators and time steps can be used in different subdomains; for a detailed discussion of Schwarz time-stepping and related machinery, the reader is referred to [24, 22]. In the numerical experiments presented herein, we restrict attention to the case where all subdomains have the same time integrator and time step, as our goal is to assess the method’s viability when coupling subdomain-local operator inference-based ROMs with each other and with subdomain-local FOMs while eliminating other confounding factors.
@@ 2.2 Operator Inference Preliminaries
Operator inference is a data-driven, non-intrusive, projection-based method for the model order reduction of PDEs, developed relatively recently by Peherstorfer and Willcox [26] as an alternative to standard Galerkin projection ROMs. It consists of two major steps:
  1. 1.
    Build a reduced basis from data.
  2. 2.
    Infer operators on the span of this reduced basis.
These two steps are described succinctly below.
@@ 2.2.1 Proper Orthogonal Decomposition
To construct a reduced basis, we will use proper orthogonal decomposition (POD) [13, 37], also known as principal component analysis (PCA) in some fields, though one might also use other methods, e.g., the reduced basis method with a greedy algorithm [30]. To perform POD, we require snapshots from some data source or FOM. Continuing the example of the heat equation from (1), a spatially discretized FOM for the heat equation will typically appear in the following form after a boundary lift:
$\dot{\mathbf{x}}=\mathbf{K}\mathbf{x}+\mathbf{B}\mathbf{g}+\mathbf{f},$ (6)
where $\mathbf{x}\in\mathbb{R}^{N}$ and $\mathbf{f}\in\mathbb{R}^{N}$ are now discretized vectors corresponding to the unconstrained state degrees of freedom (DoFs), and $\mathbf{g}\in\mathbb{R}^{m}$ discretizes the Dirichlet boundary condition. The matrix $\mathbf{K}\in\mathbb{R}^{N\times N}$ comes from discretizing the continuous Laplace operator $\Delta$, and $\mathbf{B}\in\mathbb{R}^{N\times m}$ deals with effects of the boundary term. The full state representation $\mathbf{u}\in\mathbb{R}^{N+m}$ for all DoFs is obtained by augmenting the unconstrained solution $\mathbf{x}$ with the known boundary condition $\mathbf{g}$.
After (6) has been solved in time for $\tau$ separate states inclusively between time $t=0$ and $t=T$ (i.e., $0=t_{1}<t_{2}<\dots<t_{\tau}=T$), we obtain a collection of unconstrained state snapshots,
$\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{\tau}]\in\mathbb{R}^{N\times\tau},$ (7)
[…]
$\max_{\tilde{\psi}\in H}\sum_{i=1}^{\tau}\frac{\left|\langle\mathbf{x}_{i},\tilde{\psi}\rangle\right|^{2}}{\|\tilde{\psi}\|^{2}}.$ (8)
Equation (8) is satisfied by $\psi_{1}$, the left singular vector associated with $\sigma_{1}$, the largest singular value of $\mathbf{X}$ as shown in [40]. Further basis vectors up to dimension $r\leq d$ are chosen to be orthonormal to previous basis vectors, i.e., $\langle\mathbf{\psi}_{i},\psi_{j}\rangle=\delta_{ij}$, the Kroneker delta. The columns of $\mathbf{\Psi}$ necessarily satisfy this additional property. The restriction of $\mathbf{\Psi}$ to its first $r$ columns will be referred to as $\mathbf{\Psi}_{r}$ going forwards.
[…]

### v3

@@ Abstract
This paper presents and evaluates an approach for coupling subdomain-local reduced order models (ROMs) constructed via non-intrusive operator inference (OpInf) with each other and with subdomain-local full order models (FOMs), following a domain decomposition of the spatial geometry on which a given partial differential equation (PDE) is posed. Joining subdomain-local models is accomplished using the overlapping Schwarz alternating method, a minimally-intrusive multiscale coupling technique that transforms a monolithic problem into a sequence of subdomain-local problems, which communicate through transmission boundary conditions imposed on the subdomain interfaces. After formulating the overlapping Schwarz alternating method for OpInf ROMs, we evaluate the method’s accuracy and efficiency on several test cases involving the heat equation in two spatial dimensions. We demonstrate that the method is capable of coupling arbitrary combinations of OpInf ROMs and FOMs, and that moderate speed-ups over a monolithic FOM are possible when performing OpInf ROM coupling.
@@ 1 Introduction
Despite advancements in both computer architectures and algorithms, the modeling and simulation of complex physical systems often requires tremendous computational resources. These requirements may preclude many-query analyses such as engineering design or uncertainty quantification. While projection-based reduced order models (ROMs) have shown promise to mitigate this difficulty, traditional intrusive ROMs, e.g., Galerkin [12, 36] and least squares Petrov-Galerkin (LSPG) projection ROMs [4], have their own shortcomings, including a lack of systematic refinement mechanisms, a lack of robustness, stability, and accuracy in the predictive regime, and lengthy implementation time requirements.
@@
This paper presents an approach for mitigating the aforementioned difficulties by enabling domain decomposition- (DD-)based coupling of subdomain-local ROMs with each other and/or with subdomain-local full order models (FOMs). Our approach is based on the following ingredients: (i) a decomposition of the physical domain of interest into two or more overlapping subdomains, (ii) the construction of subdomain-local ROMs and/or FOMs in each of the subdomains, and (iii) the rigorous coupling of the subdomain-local models via the Schwarz alternating method [33]. The Schwarz alternating method is based on the simple idea that, if the solution to a partial differential equation (PDE) is known in two or more regularly shaped domains, these local solutions can be used to iteratively build a solution for the union of the subdomains, with information propagating between the subdomains through carefully constructed transmission boundary conditions (BCs). We choose the Schwarz alternating method since it has a number of advantages over competing multiscale coupling methods. These advantages include its concurrent nature, its ability to couple non-conformal meshes with different element topologies and different time integrators with different time steps for dynamic problems without introducing non-physical artifacts into the solution, and its non-intrusive implementation into existing codes [24, 25].
@@
Building on our past work in developing the Schwarz alternating method as a means to couple FOMs [24, 25], intrusive projection-based ROMs [3] and physics-informed neural networks (PINNs) [37], we focus our attention herein on advancing the method to work with a non-intrusive model order reduction (MOR) technique known as operator inference (OpInf) [11, 16, 27]. Unlike traditional intrusive MOR, which requires access to the underlying FOM code in order to project the governing PDE(s) onto a reduced subspace, OpInf works by assuming a functional form (usually linear or quadratic [10]) for the ROM in terms of to-be-learned reduced operators, and solving an optimization problem offline for these operators. This procedure significantly reduces both the development time and the time-to-impact.
DD-based FOM-ROM and ROM-ROM couplings such as those proposed herein have the potential of improving the predictive viability of projection-based ROMs, by enabling the spatial localization of ROMs (via domain decomposition) and the online integration of high-fidelity information into these models (via FOM coupling). While DD-based couplings between ROMs and FOMs are not new, the majority of the literature on this topic has focused on developing intrusive coupling methods (e.g., Lagrange multipliers, optimization-based coupling) used to couple intrusive ROMs; the interested reader is referred to [3, 5, 7, 8, 13, 20, 28] and the references therein for more details. Related past work on data-driven couplings using Schwarz-like methods has focused on intrusive ROMs [6, 14, 15, 29], or on utilizing the coupling to accelerate NN training [18, 19, 37]. The proposed approach is most similar to the recent work by Farcas et al. [9], which develops a DD-based coupling of subdomain-local OpInf ROMs by learning appropriate reduced operators responsible for the coupling, and demonstrates the method on a formidable three-dimensional (3D) combustion problem. In this approach, each subdomain problem is solved once rather than by performing an iteration to convergence as done within our Schwarz framework. As a result, the subdomain-local solutions must be extended to the full domain and smoothly combined to achieve a continuous solution.
@@
The remainder of this paper is organized as follows. In Section 2, we describe the overlapping version of the Schwarz alternating method applied to our model problem, the two-dimensional (2D) unsteady heat equation. In Section 2.2, we present some OpInf preliminaries. In Section 3, we describe our software implementations of the proposed Schwarz-based coupling approach applied to OpInf ROMs, hereafter called the OpInf-Schwarz method, which makes use of the open source FEniCSx [2] and OpInf Python libraries. Numerical results are presented in Section 4. We conclude with a summary and a discussion of future work in Section 5.
@@ 2 The Schwarz Alternating Method Applied to Operator Inference ROMs
In the following sections and subsections, we consider the specific model problem of the 2D heat equation, towards addressing the challenges that are encountered in this novel combination of the Schwarz method and operator inference. We stress that neither of these techniques inherently require an assumption of linearity, noting that the authors of the original operator inference paper [27] directly contrasted their method with the linearity assumption of Dynamic Mode Decomposition (DMD) [32]. While non-linearity will likely present further challenges to overcome for the OpInf-Schwarz coupling method, the heat equation provides an initial unsteady test problem that can suggest the feasibility of the method for more complicated problems.
@@ 2.1 Schwarz Alternating Method Preliminaries
$\begin{array}[]{rcll}\dot{u}(x,t)-\Delta u(x,t)&=&0,&\text{in }\Omega\times[0,T],\\ u(x,t)&=&g(x),&\text{on }\partial\Omega\times[0,T],\\ u(x,0)&=&v(x),&\text{in }\Omega,\end{array}$ (1)
where $\Omega\in\mathbb{R}^{p}$ is an open bounded domain for $p=1,2,3$ with boundary $\partial\Omega$, $g(x)$ are is a given boundary condition function, $v(x)$ defines the initial condition, and $T>0$. Suppose we decompose the domain $\Omega$ into two overlapping subdomains $\Omega_{1}$ and $\Omega_{2}$, such that $\Omega_{1}\cup\Omega_{2}=\Omega$ and $\Omega_{1}\cap\Omega_{2}\neq\emptyset$ and, for later use, define $\overline{\Omega_{1}}$ as the closure of $\Omega_{1}$ and similarly for $\Omega_{2}$. Suppose also that we decompose the time interval $[0,T]$ into a set of non-overlapping time intervals $I_{n}=[t_{n},t_{n+1}]$, where $T\geq t_{n+1}>t_{n}\geq 0$, so that $\cup_{n}I_{n}=[0,T]$. For a given time interval $I_{n}$, the overlapping Schwarz algorithm solves the following sequence of subdomain-local problems:
$\left\{\begin{array}[]{rcll}\dot{u}_{1}^{(k+1)}-\Delta u_{1}^{(k+1)}&=&0,&\text{in }\Omega_{1}\times[t_{n},t_{n+1}]\\ u_{1}^{(k+1)}&=&g,&\text{on }(\partial\Omega\cap\overline{\Omega_{1}})\times[t_{n},t_{n+1}]\\ u_{1}^{(k+1)}&=&u_{2}^{(k)},&\text{on }\Gamma_{1}\times[t_{n},t_{n+1}],\end{array}\right.$ (2)
@@
$\left\{\begin{array}[]{rcll}\dot{u}_{2}^{(k+1)}-\Delta u_{2}^{(k+1)}&=&0,&\text{in }\Omega_{2}\times[t_{n},t_{n+1}]\\ u_{2}^{(k+1)}&=&g,&\text{on }(\partial\Omega\cap\overline{\Omega_{2}})\times[t_{n},t_{n+1}]\\ u_{2}^{(k+1)}&=&u_{1}^{(k+1)},&\text{on }\Gamma_{2}\times[t_{n},t_{n+1}],\end{array}\right.$ (3)
for Schwarz iteration $k=0,1,2,...,$ subject to initial conditions $u_{1}(x,0)=v|_{\Omega_{1}}$ and $u_{2}(x,0)=v|_{\Omega_{2}}$. In (2), $u_{i}$ for $i=1,2$ denotes the solution in subdomain $\Omega_{i}$, and $\Gamma_{i}$ is the so-called Schwarz boundary (see Figure 1). It is common to set $u_{1}^{(0)}=v|_{\partial\Omega_{1}}$ on $\Gamma_{1}$ and $u_{2}^{(0)}=v|_{\partial\Omega_{2}}$ on $\Gamma_{2}$ when initializing the Schwarz iteration process, to ensure solution compatibility with the initial condition. The iterative process in (2) and (3) continues until a set of pre-determined criteria are met. In the present work, convergence criteria are based on the Euclidean norm of the solution differences between consecutive Schwarz iterations; that is, Schwarz is deemed converged when $\epsilon_{\text{abs}}^{(k)}<\delta_{\text{abs}}$ and $\epsilon_{\text{rel}}^{(k)}<\delta_{\text{rel}}$ for some pre-specified Schwarz tolerances $\delta_{\text{abs}},\delta_{\text{rel}}>0$, where
@@
for Schwarz iteration $k=1,2,...$, until convergence.
A key advantage of the Schwarz alternating method is that it allows for the subdomains $\Omega_{1}$ and $\Omega_{2}$ to be discretized using different meshes and/or element types [24, 25]. In the case where $\Omega_{1}$ and $\Omega_{2}$ are discretized using different meshes and do not have a coincident interface, applying the Schwarz boundary condition on the Schwarz boundaries $\Gamma_{i}$ requires the construction of a projection operator; this can be done via a simple application of finite element interpolation functions readily available in most codes [23, 25].
The Schwarz iteration process (2)–(3) is converged within each time interval $[t_{n},t_{n+1}]$ before moving on to the next time interval. A key advantage of this time-stepping approach is that different time integrators and time steps can be used in different subdomains; for a detailed discussion of Schwarz time-stepping and related machinery, the reader is referred to [25, 23]. In the numerical experiments presented herein, we restrict attention to the case where all subdomains have the same time integrator and time step, as our goal is to assess the method’s viability when coupling subdomain-local operator inference-based ROMs with each other and with subdomain-local FOMs while eliminating other confounding factors.
2.1.1 FOM-FOM Schwarz Coupling
Continuing the example of the heat equation from (1), a spatially discretized monolithic FOM for the heat equation typically appears in the following form after a boundary lift:
$\dot{\mathbf{x}}=\mathbf{K}\mathbf{x}+\mathbf{B}\mathbf{g},$ (6)
where $\mathbf{x}\in\mathbb{R}^{N}$ is a discretized vector corresponding to the unconstrained state degrees of freedom (DoFs), and $\mathbf{g}\in\mathbb{R}^{m}$ discretizes the Dirichlet boundary condition. The matrix $\mathbf{K}\in\mathbb{R}^{N\times N}$ comes from discretizing the continuous Laplace operator $\Delta$, and $\mathbf{B}\in\mathbb{R}^{N\times m}$ deals with effects of the boundary condition. The full state representation $\mathbf{u}\in\mathbb{R}^{N+m}$ for all DoFs is obtained by augmenting the unconstrained solution $\mathbf{x}$ with the known boundary condition $\mathbf{g}$.
If a domain decomposition such as in Figure 1 is imposed, we may consider the subdomain-local discretized problems:
$\dot{\mathbf{x}}_{i}=\mathbf{K}_{i}\mathbf{x}_{i}+\mathbf{B}_{i}\mathbf{y}_{i},\;\;\;i=1,2,$ (7)
on $\Omega_{1}$ and $\Omega_{2}$. Let $N_{i}$ and $m_{i}$ denote the number of finite element (FE) nodes on the interior and boundary of $\Omega_{i}$, respectively. Note that the boundary of $\Omega_{i}$ includes $\Gamma_{i}$, as shown in Figure 1. We introduce the vector $\mathbf{y}_{i}\in\mathbb{R}^{m_{i}}$ to handle the subdomain-local boundary by defining $\mathbf{y}_{i}=[\mathbf{g}_{i}\;\mathbf{\gamma}_{i}]^{T}$, where $\mathbf{g}_{i}$ discretizes $g(x)$ on $(\partial\Omega\cap\overline{\Omega_{i}})$ and $\gamma_{i}$ discretizes the subdomain-local boundary condition on $\Gamma_{i}$.
The remaining quantities in (7) follow naturally from the monolithic terms in (6), with subdomain-local state representation $\mathbf{x}_{i}\in\mathbb{R}^{N_{i}}$ as well as stiffness and boundary matrices $\mathbf{K}_{i}\in\mathbb{R}^{N_{i}\times N_{i}}$ and $\mathbf{B}_{i}\in\mathbb{R}^{N_{i}\times m_{i}}$. We now use the Schwarz method to set up the subdomain-local problems as follows:
$\left\{\begin{array}[]{rcll}\dot{\mathbf{x}}_{1}^{(k+1)}&=&\mathbf{K}_{1}\mathbf{x}_{1}^{(k+1)}+\mathbf{B}_{1}\mathbf{y}_{1}^{(k+1)}\\ \gamma_{1}^{(k+1)}&=&\mathbf{x}_{2}^{(k)}\big{|}_{\Gamma_{1}},\end{array}\right.$ (8)
and
$\left\{\begin{array}[]{rcll}\dot{\mathbf{x}}_{2}^{(k+1)}&=&\mathbf{K}_{2}\mathbf{x}_{2}^{(k+1)}+\mathbf{B}_{2}\mathbf{y}_{2}^{(k+1)}\\ \gamma_{2}^{(k+1)}&=&\mathbf{x}_{1}^{(k+1)}\big{|}_{\Gamma_{2}},\end{array}\right.$ (9)
for Schwarz iteration $k=0,1,2,...,$ until convergence following (4)-(5). For work presented in this paper, we solve equations (8)-(9) using the finite element method in space and the finite difference method in time. We take the initial condition to be the interpolation of $u(x,0)=v(x)$ into the respective finite element spaces. This specific formulation is valid for conformal spatial meshes only.
@@ 2.2 Operator Inference Preliminaries
Operator Inference is a data-driven, non-intrusive, projection-based method for model order reduction, developed by Peherstorfer and Willcox [27] as an alternative to standard Galerkin projection ROMs. Similarly to other Galerkin projection ROMs, the method begins by constructing a reduced basis from data. OpInf then diverges from the standard method in the construction of the reduced operators, which are estimated using data through a regression problem.
@@ 2.2.1 Proper Orthogonal Decomposition
To construct a reduced basis, we use proper orthogonal decomposition (POD) [12, 17, 36], though one might also use other methods, e.g., the reduced basis method with a greedy algorithm [30]. To perform POD, we require snapshots from some data source or FOM.
Continuing with the heat equation example, after the monolithic problem (6) has been solved in time for $\tau$ separate states inclusively between time $t=0$ and $t=T$ (i.e., $0=t_{1}<t_{2}<\dots<t_{\tau}=T$), we obtain a collection of unconstrained state snapshots,
$\mathbf{X}=[\mathbf{x}(t_{1}),\mathbf{x}(t_{2}),\dots,\mathbf{x}(t_{\tau})]\in\mathbb{R}^{N\times\tau},$ (10)
with rank$(\mathbf{X})=d>0$. $\mathbf{X}$ yields a singular value decomposition (SVD) $\mathbf{X}=\mathbf{\Psi}\mathbf{\Sigma}\mathbf{\Phi}^{*}$, where $\mathbf{\Psi}\in\mathbb{R}^{N\times N}$ is an orthogonal matrix whose first $d$ columns form a basis for the column space of $\mathbf{X}$. The columns of $\mathbf{\Psi}$, $\mathbf{\psi}_{i}$ for $i=1,\dots r$ with $r\leq d$ form an optimal $r$-dimensional basis for the columns of $\mathbf{X}$ in an $\ell^{2}$ sense. The restriction of $\mathbf{\Psi}$ to its first $r$ columns is referred to as $\mathbf{\Psi}_{r}\in\mathbb{R}^{N\times r}$ going forwards.